Please solve with steps

A sticky projectile collides horizontally with a stationary block on a horizontal surface. The collision is perfectly inelastic and the block slides 7.50 m before stopping. The mass of the projectile is 12.0 g, the blocks mass is 100 g, and the coefficient of friction between the block and the surface is 0.65. What was the projectiles speed before impact? (15 Points)
Answer #1

deacceleration = ug

where u = coefficient of friction

g= acceleration due to gravity

= 0.65*9.8 = 6.37 m/s2

using the equation of linear motion we have

v^2 - u^2 = 2aS

0 - u^2 = 2*-6.37*7.5

u = 9.77 m/s

speed when they both stick together is 9.77 m/s

let the speed of projectile before impact be v

by the conservation of momentum we have

m1u1 + m2u2 = (m1 + m2)V

100*0 + 12*v = (100 + 12)*9.77

v = 91.23 m/s